OpenStudy (anonymous):

If x,y are the integral sollutions of 2x^2 -3xy - 2y^2 = 7 , then mod(x+y) =

ganeshie8 (ganeshie8):

I would start by trying to factor the left quadratic

OpenStudy (anonymous):

How to do that

ganeshie8 (ganeshie8):

its just a quadratic, split the middle term

ganeshie8 (ganeshie8):

-4+1 = -3 -4*1 = -4

ganeshie8 (ganeshie8):

2x^2 -3xy - 2y^2 = 7 2x^2 - 4xy - xy - 2y^2 = 7 group terms and factor

OpenStudy (anonymous):

was doing my net disconnected

OpenStudy (anonymous):

2x(x-2y)-y(x-2y) (2x-y)(x-2y)

ganeshie8 (ganeshie8):

2x^2 - 4xy + xy - 2y^2 = 7 2x(x-2y) + y(x-2y) = 7 (x-2y)(2x+y) = 7

ganeshie8 (ganeshie8):

the only factors that multiply to 7 are : {1, 7} or {-1, -7}

OpenStudy (anonymous):

ya sry

OpenStudy (anonymous):

i think i got the next step

OpenStudy (anonymous):